A boy made 40 quarts of lemonade, which contained 20% lemon. How much water
would he have to add to reduce it to 10% lemon?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

A boy made 40 quarts of lemonade, which contained 20% lemon. How much water
would he have to add to reduce it to 10% lemon?
-----------------------------
Equation:
lemon + lemon = lemon
0.20*40 + 0*x = 0.10(40+x)
Multiply thru by 100 to get:
20*40 + 0 = 10*40 + 10x
10x = 10*40
x = 40 guarts (amt. of water needed for the mixture)
==================
Cheers,
Stan H.

Answer by richwmiller(17219)

.20*40+0*q=.10(40+q)
q=40
which makes sense since we want half the potency so we double the quantity